# Decom-Renorm-Merge: Model Merging on the Right Space Improves Multitasking

### v1

@@ Appendix D Theoretical Bound of Difference Between Individual Task and the Concatenated Matrix
$\sigma_{i}(M)=\sqrt{k}\sigma_{i}(A),$
@@
Building on the previous arguments, we hereby establish the bound:
Proposition 2.
Suppose that $A$ is an arbitrary weight delta $\Delta W_{t}$, and $M\in\mathbb{R}^{m\times kn}$ is a concatenation of $k$ copies of A, where $k$ is the total number of tasks and $r$ is the rank of $M$. Let $\widetilde{M}=[\Delta W_{1}\;\dots\;\Delta W_{k}]$ be the concatenation of the collection $\{\Delta W_{j}\}_{j=1}^{k}$. We define the perturbation matrix
$E=\widetilde{M}-M=[E_{1}\;\dots E_{k}],\quad\text{with $E_{j}=\Delta W_{j}-A$ \quad for each $j\in[1,k]$.}$
Then, the difference between $i$-th singular value of concatenated matrix $\widetilde{M}$ and individual $\Delta W_{t}$ for all $i\in[1,r]$ and $t\in[1,k]$ is bounded by
$|\sigma_{i}(\widetilde{M})-\sqrt{k}\sigma_{i}(A)|\leq\frac{1}{\sqrt{k}\sigma_{i}(A)}\sum_{j=1}^{k}\left(2\|A\|_{2}\|E_{j}\|_{2}+\|E_{j}\|_{2}^{2}\right).$
@@
Then, we substitute $\sigma_{i}(M)$ with $\sqrt{k}\sigma_{i}(A)$, following Proposition 1:
@@
This result quantifies the bounded behavior of change to each individual task’s weight delta, such that the alteration induced by horizontal concatenation of weight deltas will never exceed a specific amount described by the bound. From this derived inequality, it is evident that larger singular values $\sigma_{i}(\Delta W_{t})$ will exhibit a tighter bound due to the reciprocal term $\frac{1}{\sigma_{i}(\Delta W_{t})}$ on the right hand side. The coefficient $\sqrt{k}$ naturally arises for our DRM approach, which involves concatenating $k$ matrices; however, this factor is generally not problematic as the number of tasks $k$ is typically small and remains relatively consistent for each merging setting. Likewise, vertically concatenated matrices can also be bounded by going through the same steps on the transposed version, producing a very similar inequality due to the invariance under transpose property of singular value and spectral norm. We then discuss the indication of this singular value bound:
Bounded behavior of weight deltas concatenation.
Intuitively, when comparing weight deltas of different models during interference mitigation, it is essential to preserve the top principle features of each model to ensure a fair comparison across tasks; otherwise, certain tasks could be degraded unknowingly. Our derived bound addresses this by indicating a bounded degree of change, especially for the important feature bases characterized by high singular values. This bound, inherent to the properties of the concatenation process, assures the preservation of salient information from each model, facilitating interference reduction and merging in the subsequent steps.
@@ E.1 Singular Value Decomposition
Different orders of weight delta concatenation: $[A\;B]$ and $[B\;A]$ should produce the same shared space represented by $U$ (or very similar in practice). This invariance is crucial in ensuring the robustness of the decomposition process. As we treat each task as equal peers, shuffling the task concatenation order should not generate different merging outcomes.
@@ Appendix F Elaboration of Sign Agreement Analysis
Note that that the minimum sign agreement is $50\%$, when there are an equal numbers of positive and negative entries. Notably, for the bin of $0.9-1.0$ sign agreement level, transforming back to the original space showcases a sign agreement drop of approximately 3 folds, on both models we experimented on. This supports our conclusion in Section 6.2.
@@ H.1 Effect of rank truncation
Figure 6: Performance of T5-Base and ViT-B/32 when dropping small singular values. A rank drop rate of $k$ means retaining only the top $(1-k)$ portion of the singular values before renormalization (multiplying vector norms back into the singular values during renormalization could cause changes in singular value ordering). We see that the performance remains relatively stable, with only a slight decreasing trend. The performance only begins to degrade noticeably after $k$ is greater than 0.8.
@@ I.3.2 Datasets
We evaluated our approach on the following datasets from the GLUE benchmark [78]: QNLI [^13] https://huggingface.co/datasets/nyu-mll/glue/viewer/qnli, MNLI [^14] https://huggingface.co/datasets/nyu-mll/glue/viewer/mnli, RTE [^15] https://huggingface.co/datasets/nyu-mll/glue/viewer/rte, SST-2161616https://huggingface.co/datasets/nyu-mll/glue/viewer/sst2, and CoLA [^17] https://huggingface.co/datasets/nyu-mll/glue/viewer/cola, chosen to cover a range of natural language understanding tasks. Since every dataset here lacks publicly available test labels, we used the original validation sets for testing, and created new validation splits by reserving $10\%$ of the training data, except for SST-2 and CoLA, which use only 5%, using stratified sampling with a fixed random seed of 0. The dataset statistics are summarized in Table 9.

### v2

@@ D.2 Bound of Difference Between Individual Task and the Concatenated Matrix
$\sigma_{i}(M)=\sqrt{k}\sigma_{i}(A)$
@@
Building on the previous arguments, we now establish the proof of the main Proposition 2 stated in the Section 6.3:
@@
Then, we substitute $\sigma_{i}(M)$ with $\sqrt{k}\sigma_{i}(A)$, following Lemma 5:
@@
The coefficient $\sqrt{k}$ naturally arises for our DRM approach, which involves concatenating $k$ matrices; however, this factor is generally not problematic as the number of tasks $k$ is typically small and remains relatively consistent for each merging setting. Likewise, vertically concatenated matrices can also be bounded by going through the same steps on the transposed version, producing a very similar inequality due to the invariance under transpose property of singular value and spectral norm.
Limitation.
While Proposition 2 establishes the upper bound of difference between $i$-th singular values of (1) weight delta of an individual task, and (2) jointly concatenated weight delta; however, note that $i$-th basis vectors of the two matrices doesn’t necessarily correspond to the exact same basis direction.
@@ E.1 Singular Value Decomposition
Different orders of weight delta concatenation: [AB][A\;B] and [BA][B\;A] should produce the same shared space represented by $U$ (or very similar in practice). This invariance is crucial in ensuring the robustness of the decomposition process. As we treat each task as equal peers, shuffling the task concatenation order should not generate different merging outcomes.
@@ Appendix F Elaboration of Sign Agreement Analysis
Note that that the minimum sign agreement is $50\%$, when there are an equal numbers of positive and negative entries. Notably, for the bin of 0.91.00.9-1.0 sign agreement level, transforming back to the original space showcases a sign agreement drop of approximately 3 folds, on both models we experimented on. This supports our conclusion in Section 6.2.
@@ H.1 Effect of Rank Truncation
Figure 6: Performance of T5-Base and ViT-B/32 when dropping small singular values. A rank drop rate of $k$ means retaining only the top (1k)(1-k) portion of the singular values before renormalization (multiplying vector norms back into the singular values during renormalization could cause changes in singular value ordering). We see that the performance remains relatively stable, with only a slight decreasing trend. The performance only begins to degrade noticeably after $k$ is greater than 0.8.
@@ I.3.2 Datasets
We evaluated our approach on the following datasets from the GLUE benchmark [78]: QNLI [^14] https://huggingface.co/datasets/nyu-mll/glue/viewer/qnli, MNLI [^15] https://huggingface.co/datasets/nyu-mll/glue/viewer/mnli, RTE [^16] https://huggingface.co/datasets/nyu-mll/glue/viewer/rte, SST-2171717https://huggingface.co/datasets/nyu-mll/glue/viewer/sst2, and CoLA [^18] https://huggingface.co/datasets/nyu-mll/glue/viewer/cola, chosen to cover a range of natural language understanding tasks. Since every dataset here lacks publicly available test labels, we used the original validation sets for testing, and created new validation splits by reserving $10\%$ of the training data, except for SST-2 and CoLA, which use only 5%, using stratified sampling with a fixed random seed of 0. The dataset statistics are summarized in Table 9.